I think I did the first part correctly, but got stuck on what to do next..can you please help me? Thank you!
If 7x -3 = 25, then what is 2x + 9

3+3= 0
25 + 3 = 28
7x = 28
7x divided by 7 = x
28 divided by 7 = 4
Is that right so far, and what do I do next if it is?
Thank you in advance!!!!!

Answer by tutorcecilia(2152)

If 7x -3 = 25, then what is 2x + 9
7x-3=25 [solve for x]
7x-3+3=25+3
7x=28
7x/7=28/7
x=4
.
.
Plug (x=4) into the second equation:
2x+9
2(4)+9 [Simplify]
8+9
17 [Final answer for second problem